OpenStudy (anonymous):

show that \( 1+\frac{1}{4}+\frac{1}{9}+\frac{1}{16}+...=\frac{1}{6} \pi^2 \)

OpenStudy (perl):

this is famous basel problem solved by euler i believe

OpenStudy (anonymous):

aha then ?

OpenStudy (perl):

there are many proofs online, wikipedia

OpenStudy (anonymous):

mmm

OpenStudy (perl):

http://en.wikipedia.org/wiki/Basel_problem

OpenStudy (perl):

The Basel problem asks for the exact sum of this series (in closed form), as well as a proof that this sum is correct. Euler found the exact sum to be π2/6 and announced this discovery in 1735. His arguments were based on manipulations that were not justified at the time, and it was not until 1741 that he was able to produce a truly rigorous proof.

OpenStudy (anonymous):

im given this Qn , and it asked to make usufull of fourier to show this

OpenStudy (perl):

what Qn are you given, sorry i dont understand your english

OpenStudy (anonymous):

you must write fourier series of \(x^2\) for some symmetric interval

OpenStudy (anonymous):

use the problem " find the fourier series of the function f(x) with period p=2l f(x)=x^2 (-1<x<1) p=2"

OpenStudy (anonymous):

ohk cool @mukushla can u help me with that ?

OpenStudy (anonymous):

It's a well known problem, yes, let's work it together if you want

OpenStudy (anonymous):

yeah sure ^_^

OpenStudy (anonymous):

i tried on x^2 ill type the final i got hope its right

OpenStudy (anonymous):

\(\large \frac{\pi^2}{3}+\sum_{n=1}^{\infty} a_n \cos nx =\frac{\pi^2}{3} +a_1\cos x+a_2\cos2 x ++a_3\cos3 x\) =\(\large \frac{\pi^2}{3} -4\cos x+\cos2 x -\frac{4}{9}\cos 3x+\frac{1}{4}\cos 4x+...\)

OpenStudy (anonymous):

Can you provide general form of \(a_n\) ? :-)

OpenStudy (anonymous):

yess a_0=\(\pi^2/3\) ill type a_n

OpenStudy (anonymous):

\(4\frac{(-1)^n}{n^2}\)?

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

well, now let \(x=0\), see what happens

OpenStudy (anonymous):

for the series ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\large \frac{\pi^2}{3} -4 +1 -\frac{4}{9} +\frac{1}{4}-\frac{4}{25} +... \)

OpenStudy (anonymous):

so\[0=\frac{\pi^2}{3}+\sum_{n=1}^{\infty} a_n \]\[\sum_{n=1}^{\infty} \frac{4(-1)^n}{n^2}=-\frac{\pi^2}{3}\]\[\sum_{n=1}^{\infty} \frac{(-1)^n}{n^2}=-\frac{\pi^2}{12}\]so we need to plug in another number for \(x\)

OpenStudy (anonymous):

aha ic..

OpenStudy (anonymous):

what do you think? what makes that \((-1)^n\) disappear?

OpenStudy (perl):

this is fourier series, sorry i have not learned this :o

OpenStudy (anonymous):

mmm pi ?

OpenStudy (anonymous):

its ok perl :D

OpenStudy (anonymous):

\(\pi\) is not on our interval

OpenStudy (anonymous):

the interval is given for pre Qn

OpenStudy (anonymous):

I think there is a little mistake in your final answer, that must be\[x^2=\frac{\pi^2}{3}+\sum_{n=0}^{\infty} \frac{4(-1)^n}{n^2} \cos n\color\red{\pi} x\]

OpenStudy (anonymous):

mmm why ?

OpenStudy (anonymous):

cuz of the interval ?

OpenStudy (anonymous):

i solved this for -pi<x<pi

OpenStudy (anonymous):

took period of 2 pi

OpenStudy (anonymous):

aha i thought u used \((-1,1)\)

OpenStudy (anonymous):

no ;)

OpenStudy (anonymous):

so you are quite right :-) you know why? \[x^2=\frac{\pi^2}{3}+\sum_{n=1}^{\infty} \frac{4(-1)^n}{n^2} \cos n x\]let \(x=\pi\) \[\pi^2=\frac{\pi^2}{3}+\sum_{n=1}^{\infty} \frac{4(-1)^n}{n^2} (-1)^{n}\]\[\frac{\pi^2}{6}=\sum_{n=1}^{\infty} \frac{1}{n^2}\]

OpenStudy (anonymous):

do u think i should make the interval -1<x<1 to have n pi x ?

OpenStudy (anonymous):

ohh lol i forget totake x^2 xD ok i got it nww tyyyyyy <3

OpenStudy (anonymous):

very welcome <3

OpenStudy (perl):

is there an introduction to fourier series online?

OpenStudy (anonymous):

sure, just google it ;-)

OpenStudy (perl):

that deserves a medal

OpenStudy (anonymous):

:D